OpenStudy (anonymous):

How many pounds of an $83 per pound substance must be mixed with an $11 per pound substance to get 240 lbs of an alloy which will cost $28 per pound? Round to the nearest pound. 57 lbs 183 lbs 205 lbs 35 lbs

OpenStudy (anonymous):

make a chart

OpenStudy (anonymous):

You can first find the % of each substance needed expressed as a decimal. The amount of the first substance can be x. The amount of the other substance then will be 1-x. Set up an equation as follows: 83x+11(1-x)=28 83x+11-11x=28 72x=17 x=17/72 which is about 0.24 In other words, you need about 24% of the first substance. 0.24*240=57.6, so the closest answer you have is 57 lbs. Cheers!